7. How large was $16.5 trillion. The exchange rate in 2014 was 61.0 rupees per dollar. India turns out to have lower prices than the United States (this is true more generally for poor countries): the price level in India (converted to dollars) divided by the price level in the United States was 0.280 in 2014. (a) What is the ratio of Indian GDP to U.S. GDP if we dont take into account the differences in relative prices and simply use the exchange rate to make the conversion? (b) What is the ratio of real GDP in India to real GDP in the United States in common prices? (c) Why are these two numbers different?
Answer #1

(a) If we simply use the exchange rate, i.e. 61 R/D, Indian GDP= $ 119 Tr/ 61 and US GDP =$16.5 Tr, this gives the ratio of
Indian GDP/ US GDP= (119/61) / 16.5 = 0.12 i.e. Indian economy is ~12% the size of US economy.


(b) Real GDP in India/ Real GDP in US = Nominal Indian GDP/ P (India) / Nominal US GDP/ P (US)
= Nominal Indian GDP/ Nominal US GDP* (P(India)/ P(US)) = 0.12/0.280 = 0.43
Hence, in terms of common prices, Indian economy is 43% of the US economy.

(c) This is because of the difference in price levels in the two countries, which is not correctly reflected in the country's exchange rate. This may be due to: India systematically undervaluing its Rupee or the general subsidised prices prevailing in the Indian market or the excessive Demand for US Dollar, as it is seen as an international currency and used by many developing countries, pushing up its demand and therefore its value!